OpenStudy (anonymous):

In the sequence of non-zero numbers x^1, x^2, x^3, x^4,..., the nth term is x^n. If, from the third term on, each term is the sum of the preceding two terms, what are all possible values of x?

OpenStudy (anonymous):

\[x^3=x^2+x\] i guess is what it is saying

OpenStudy (anonymous):

and apparently \(x\neq 0\) so you can solve \[x^3-x^2-x=0\] \[x(x^2-x-1)=0\] solve \[x^2-x-1=0\] use quadratic formula

OpenStudy (anonymous):

you good with this or need more help?

OpenStudy (anonymous):

I'm... still a little bit confused, I think.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

If, from the third term on, each term is the sum of the preceding two terms

OpenStudy (anonymous):

this means the third term, which is \(x^3\) is the sum of the previous two terms, \(x\) and \(x^2\)

OpenStudy (anonymous):

so we can write the equation \[x^3=x^2+x\]

OpenStudy (anonymous):

ok, that makes sense now...

OpenStudy (anonymous):

wait a second...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i just read back over the first part... quadratic formula...

OpenStudy (anonymous):

once you have the equation \[x^3-x^2-x=0\] you can factor as \[x(x^2-x-1)=0\]

OpenStudy (anonymous):

one solution is \(x=0\) but you are told at the beginning that \(x\) is not zero, so you can now solve \[x^2-x-1=0\]

OpenStudy (anonymous):

right... I think i have an answer, but i'm not entirely sure how to type it... one plus-minus root five, all over two?

OpenStudy (anonymous):

Or did I just do something completely different?

OpenStudy (anonymous):

yes that is the answer

OpenStudy (anonymous):

Ok, thanks! ^_^

OpenStudy (anonymous):

it might be worth checking, but is should be right

OpenStudy (anonymous):

btw, \(\frac{1+\sqrt{5}}{2}\) is a famous number, sometime denoted as \(\phi\) or the "golden ratio" if you google it you will see lots

OpenStudy (anonymous):

Huh. I'll remember that. It does look kind of familiar

OpenStudy (anonymous):

it is related to fibonacci numbers

OpenStudy (anonymous):

If k is real, what is the only real number that could be a multiple root of x^3 +kx+1=0?

OpenStudy (anonymous):

oh now i have to think

OpenStudy (anonymous):

we can do this the long way, i am not sure there is a simple way

OpenStudy (anonymous):

Ok, so now I know the answer I would usually put down in this situation is wrong. So how does the long way work?

OpenStudy (anonymous):

expand \((x-a)^2(x-b)\) if the two roots are say \(a\) and \(b\) and then equate

OpenStudy (anonymous):

you get \[(x-a)^2(x-b)=x^3-(2a+b)x^2+(a^2+2ab)x-a^2b\] if my algebra is right

OpenStudy (anonymous):

and since there is no \(x^2\) term this tells you \(2a+b=0\)

OpenStudy (anonymous):

you also know \(-a^2 b=1\)

OpenStudy (anonymous):

Whoa. How did you get that? The part right after you've expanded? I got something totally different. Or maybe I'm just working in circles and getting nowhere... I don't know what i just did. One second...

OpenStudy (anonymous):

you have to grind it out and then combine like terms, but i cheated

OpenStudy (anonymous):

So, there's nothing in here that says x^3-2ax^2+a^2x-bx^2+2abx-a^2b=0? Or did you get that before combining like terms?

OpenStudy (anonymous):

Wait , are there even like terms in here? I think I just blew this way out of proportion...

OpenStudy (anonymous):

There's nothing to go with x^3...

OpenStudy (anonymous):

yes you are right, on the right track for sure

OpenStudy (anonymous):

you got this \[ x^3-2ax^2+a^2x-bx^2+2abx-a^2b\] right?

OpenStudy (anonymous):

but what about -2ax^2 and -bx^2? they both have x^2, but would the coefficient be -2ab? and yes, i got that=0

OpenStudy (anonymous):

\(-2ax^2-bx^2=(-2a-b)x^2=-(2a+b)x^2\) is the algebra i used to get the coefficient of the \(x^2\) term

OpenStudy (anonymous):

oh like solving a literal equation, kind of?

OpenStudy (anonymous):

oh, i see what you did, no you don't multiply to get \(-2abx^2\) you add

OpenStudy (anonymous):

yeah kind of, but really it is "factoring" as in "factor the \(x^2\) out of the expression \[-2ax^2-bx^2\]

OpenStudy (anonymous):

so i multiplied instead of added, and i didnt even multiply correctly, now that i look at it. oh dear.

OpenStudy (anonymous):

similarly for the \(x\) term you get \((a^2+2ab)x\) but that is not really important here

OpenStudy (anonymous):

hope it is clear what the real problem is, find the polynomial that has no \(x^2\) term

OpenStudy (anonymous):

so, now what about a^2x and -a^2b? that would be... (x-b)a^2, right?

OpenStudy (anonymous):

no, you want all in terms of \(x\), not \(a^2\) we just made up the \(a\) as a root

OpenStudy (anonymous):

the like terms are in terms of \(x\)

OpenStudy (anonymous):

Oh. Lord, Okay. so what would I do with those two terms, then? aren't they like terms?

OpenStudy (anonymous):

and \(-a^2b\) is the constants, because there is no \(x\) in it, so we know \(-a^2b=1\)

OpenStudy (anonymous):

wait. the like terms are in terms of x. ok. hold on...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so do i need to solve x^3-(2a+b)x^2+1=0?

OpenStudy (anonymous):

if -a^2b=1?

OpenStudy (anonymous):

lets back up a second

OpenStudy (anonymous):

good idea.

OpenStudy (anonymous):

when you combine like terms (in \(x\)\) you should get \[(x-a)^2(x-b)=x^3-(2a+b)x^2+(a^2+2ab)x-a^2b\]

OpenStudy (anonymous):

and this must be equal to \(x^3+kx+1\)

OpenStudy (anonymous):

ok, i see that first part now...

OpenStudy (anonymous):

the important point is this: \(x^3+kx+1\) has no \(x^2\) term in it

OpenStudy (anonymous):

right. so... do i set the two expressions equal to each other?

OpenStudy (anonymous):

forget about the \(k\) we don't need it what has to be true is that \(2a+b=0\)

OpenStudy (anonymous):

because there is no \(x^2\) term

OpenStudy (anonymous):

and also you know the constant is \(-a^2b=1\)

OpenStudy (anonymous):

i mean x^3+kx+1=x^3-(2a+b)x^2+(a^2+2ab)x-a^2b? ok, back up. im lost again.

OpenStudy (anonymous):

yes you know \[ x^3+kx+1=x^3-(2a+b)x^2+(a^2+2ab)x-a^2b\]

OpenStudy (anonymous):

so they are equal?

OpenStudy (anonymous):

they must be equal, yes, because you are given at the outset that the polynomial is \(x^3+kx+1\)

OpenStudy (anonymous):

so now lets concentrate on finding the roots, which are \(a\) and \(b\)

OpenStudy (anonymous):

ok. so... how do i do that? unless i just square root it, but that leaves me with a bunch of radicals, right?

OpenStudy (anonymous):

now it the time to equate like coefficents

OpenStudy (anonymous):

the coefficient of the \(x^2\) term in \(x^3+kx+1\) is \(0\) and therefore \[2a+b=0\]

OpenStudy (anonymous):

ok... so... let me think...

OpenStudy (anonymous):

wait, why 0?

OpenStudy (anonymous):

\[x^3+kx+1=x^3-(2a+b)x^2+(a^2+2ab)x-a^2b\]

OpenStudy (anonymous):

oh x^2. not x^3. got it

OpenStudy (anonymous):

we also see that \(-a^2b=1\)

OpenStudy (anonymous):

how do we see that?

OpenStudy (anonymous):

look above

OpenStudy (anonymous):

oh. OH. i think im understanding a little better.

OpenStudy (anonymous):

if you plugged in 0 for x and solved, that'd be it, right? 1=-a^2b?

OpenStudy (anonymous):

\[x^3+kx+1=x^3-\overbrace{(2a+b)}^0x^2+\overbrace{(a^2+2ab)}^kx\overbrace{-a^2b}^1\]

OpenStudy (anonymous):

yes, that is another way to see it!

OpenStudy (anonymous):

aha! makes sense now.

OpenStudy (anonymous):

so now we can solve \[2a+b=0\] \[-a^2b=1\] for \(a\) and \(b\)

OpenStudy (anonymous):

so now my equation looks like x^3+kx+1=x^3+(a^2+2ab)x+1, right? and i solve that?

OpenStudy (anonymous):

we need the roots, that is, we need \(a\) and \(b\)

OpenStudy (anonymous):

ok, so im still not sure how to do that without ending up with a bunch of radicals in the equation. or can i just get rid of those easily enough after i find the roots? no, bc then i'd be going in a circle...

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

now it is just algebra, if \(2a+b=0\) then \(b=-2a\) and if \(-a^2b=1\) if you replace \(b\) by \(-2a\) you get \[-a^2(-2a)=1\\ 2a^3=1\\ a=\frac{1}{\sqrt[3]{2}}\]

OpenStudy (anonymous):

at least that is what i got,

OpenStudy (anonymous):

oh. just algebra. my bad; i over-thought again. ok, so... once you get there, you'd multiply by the denominator, right?

OpenStudy (anonymous):

lets go back up and read the question

OpenStudy (anonymous):

took a while it says this "If k is real, what is the only real number that could be a multiple root of x^3 +kx+1=0"

OpenStudy (anonymous):

ugh. what have i done? i thought that's what you did to get a radical out of the denominator...?

OpenStudy (anonymous):

oh if you want, sure go ahead i just wanted to make sure we knew we had the answer

OpenStudy (anonymous):

what?!

OpenStudy (anonymous):

i think we found it, though, it should be \[x=\frac{1}{\sqrt[3]{2}}\]